please help me solve this word problem:
the second of two numbers is 5 more than twice the first. their sum is 80. find the numbers.

Answer by checkley77(12844)

x+y=80 or y=80-x
y=2x+5
2x+5=80-x
2x+x=80-5
3x=75
x=75/3
x=25 ans. for the smaller number.
y=80-25
y=55 ans. for the larger number.